# AnoVox: A Benchmark for Multimodal Anomaly Detection in Autonomous Driving

### v1

@@ 1 Introduction
    First benchmark to include content and temporal anomalies, providing multimodal sensor data, ego-vehicle state information, and ground truth in all modalities as well as a spatial voxel representation
@@ 2 Related Work
Dataset Sensors Size
@@
FS Lost and Found Cam. 375 Recording Content 1 Sem. mask
FS Static Cam. 1,030 Data Augmentation Content 1 Sem. mask Cityscapes
@@
StreetHazards Cam. 1,500 Simulation Content 1 Sem. mask CARLA
BDD-Anomaly Cam. 810 Class Exclusion Content 3 Sem. mask BDD100K
@@
RoadAnomaly21 Cam. 110 Web Sourcing Content 1 Sem. mask Cityscapes
RoadObstacle21 Cam. 412 Recording Content 1 Sem. mask Cityscapes
@@
Recording Content 13 Inst. mask Cityscapes
CARLA-WildLife Cam., Depth 1,210 Simulation Content 18 Inst. mask CARLA
@@
Lost and Found [59] Stereo Cam. 2,104 Recording Content 42 Sem. mask
WD-Pascal [6] Cam. 70 Data Augmentation Content 1 Sem. mask WildDash
Vistas-NP [35] Cam. 11,167 Class Exclusion Content 4 Sem. mask Mapillary Vistas
MUAD [32] Cam., Depth 4,641 Sim., Class Exclusion Content 9 Sem. mask MUAD
@@
In autonomous driving, anomaly detection has been classically evaluated on a small set of benchmarks. For our analysis of related works, we have included works with published open-access perception datasets from an ego perspective that provide pixel- or point-wise ground truth. Thus, we neglected frameworks that do not provide explicit data [74, 65, 37, 73, 62] and such that provide only frame-wise annotations [64, 76, 67, 2]. For our overview of related works as shown in Table 1, we updated and extended the survey of Bogdoll et al. [15], comparing now 22 datasets from the literature.
Figure 2: Overview: With Anovox, highly configurable scenarios can be created that either represent normality or include content or temporal anomalies. Generated datasets include rich labels and ground truth in, 2D, 3D and a spatial voxel space.
@@
Finally, most benchmarks do not provide a definition of normality [9, 23, 40]. This makes it especially hard for un- and self-supervised methods to precisely detect anomalies if the semantic training distribution is not fully known. Especially desirable is a well-defined normality that allows for the generation of compliant training data. This is only possible in simulation, where full control over both the training and evaluation data is available. Sometimes, even unlabeled anomalies occur in evaluation data [69].

### v2

@@ 1 Introduction
    First benchmark to provide content and temporal anomalies, adaptable multimodal sensor data, ego-vehicle state information, and ground truth in all modalities, as well as a spatial voxel representation
@@ 2 Related Work
Dataset Data Size
@@
FS Lost and Found Cam. 375 Recording Content 1 Sem. mask (2D)
FS Static Cam. 1,030 Data Augmentation Content 1 Sem. mask (2D) Cityscapes
@@
StreetHazards Cam. 1,500 Simulation Content 1 Sem. mask (2D) CARLA
BDD-Anomaly Cam. 810 Class Exclusion Content 3 Sem. mask (2D) BDD100K
@@
RoadAnomaly21 Cam. 110 Web Sourcing Content 1 Sem. mask (2D) Cityscapes
RoadObstacle21 Cam. 412 Recording Content 1 Sem. mask (2D) Cityscapes
@@
Recording Content 13 Inst. mask (2D) Cityscapes
CARLA-WildLife Cam., Depth 1,210 Simulation Content 18 Inst. mask (2D) CARLA
@@
Lost and Found [62] Stereo Cam. 2,104 Recording Content 42 Sem. mask (2D)
WD-Pascal [6] Cam. 70 Data Augmentation Content 1 Sem. mask (2D) WildDash
Vistas-NP [36] Cam. 11,167 Class Exclusion Content 4 Sem. mask (2D) Mapillary Vistas
MUAD [32] Cam., Depth 4,641 Sim., Class Exclusion Content 9 Sem. mask (2D) MUAD
@@
In autonomous driving, anomaly detection has been classically evaluated on a small set of benchmarks. For our analysis of related works, we have included works with published open-access perception datasets from an ego perspective that provide pixel- or point-wise ground truth. Thus, we neglected frameworks that do not provide explicit data [77, 68, 38, 76, 65], such with missing or incomplete data [34, 52, 48], and any that provide only frame-wise annotations [67, 79, 70, 2]. For our overview of related works as shown in Table 1, we updated and extended the survey of Bogdoll et al. [15], comparing now 22 datasets from the literature.
Figure 2: Overview: With Anovox, highly configurable scenarios can be created that either represent normality or include content or temporal anomalies. Generated datasets include rich labels and ground truth in 2D, 3D, and a spatial voxel space.
@@
Most benchmarks do not provide a definition of normality [9, 23, 41]. This makes it especially hard for un- and self-supervised methods to precisely detect anomalies if the semantic training distribution is not fully known. Especially desirable is a well-defined normality that allows for the generation of compliant training data. This is only possible in simulation, where full control over both the training and evaluation data is available. Sometimes, even unlabeled anomalies occur in evaluation data [72].
The Synthetic Dataset for Anomaly and Corner Case Detection (SDAC) by Gong et al. [34] was developed concurrently to AnoVox and shows the highest similarities to our work. They provide multimodal camera and lidar data from the CARLA environment. Labels are provided in the form of 2D and 3D bounding boxes as well as panoptic segmentation masks.
They provide a small and fixed dataset representing normality. However, their definition is very narrow and only consists of "clear, sunny weather" scenarios with four normal classes: car, motorcycle, bicycle, and person [34].
They provide content and temporal anomalies. Their content anomalies in the "TruckBus" set are no true anomalies, but the withheld classes truck and bus. The "Props" subset includes small objects, like traffic cones, or boxes. The "Accident" subset combines unknown weather conditions with known objects in atypical positions, such as overturned vehicles or pedestrians who are lying down. In all cases, the anomalies are spawned within the drivable area of the frames.
In addition, they also provide scene anomalies that include known objects with weather shifts. However, as we do not focus on frame-wise labels, they are irrelevant here.
Their scenario anomalies focus on "pre-crash and crash scenarios" [34]. They provide seven manually generated scenario types, also including sudden breaking scenarios.
"Download link for SDAC-Scenario: [Comming Soon]" –> ignore because unpublished!